OpenStudy (anonymous):

-4(5x-6y)-(4x-9) simplify

OpenStudy (anonymous):

have any answer choices

OpenStudy (anonymous):

You have to distribute the negative four outside of the parentheses. So you have to take \[-4\times5x + -4\times-6y\]

OpenStudy (anonymous):

For the other parentheses, you have to multiply by -1, because when there no number outside, it means that there is only one of them, so \[4x \times -1 and -1 \times -9\]

OpenStudy (anonymous):

Then you add them all up.